OpenStudy (anonymous):

Simplify 2u^5w^-1*4v^-8u^5*4v^8w

OpenStudy (zzr0ck3r):

can you use brackets

OpenStudy (callisto):

I assume your question to be:\[2u^5w^{-1} \times 4v^{-8}u^5\times 4v^8w\]\[=(2\times 4 \times 4)u^{5+5}w^{-1+1} v^{-8+8}\]\[=...\]